OpenStudy (anonymous):

8 4/9 multiplied by 1/5. Change mixed number to improper.

OpenStudy (saifoo.khan):

\[\frac{76}{45}\]

OpenStudy (anonymous):

How do you change a mixed to improper? We'll say 8=a, 4=b, and 9=c in order to get an improper, you take \[a*c+b \over c\] so that is 72+4 = 76 \[76 \over 9\] times that by \[1 \over 5\] and you get \[76 \over 9*5\] \[76 \over 45\]

OpenStudy (anonymous):

(8x9+4)/9 x 1/5 76/9 x 1/5 = 76/45 or 1 31/45